A uniform electric field E = - 1000 N/C i is shown9
Answer #1

(a) Since AB is perpendicular to the direction of electric field, A and B are equipotential points. So, VB - VA = 0

(b) VC - VB = E * dCB = -1000 * (4 * 1 * 10-2) = -40 V

(c) Since VA = VB, we have,

VC - VA = VC - VB = -40 V